1) the diagonal of a square is 17m, what is the side length?
2) the perimeter of an equilateral triangle is 27 inches. What is the area?
3)a tree is 15 ft tall casts a shadow which forms an angle of 34 digrees with the ground. how long is the shadow to the nearest tenth?

Answer by edjones(8007)

1)Let the side length = s , hypotenuse = c
2s^2=c^2 Pythagoras
2s^2=17^2
2s^2=289
s^2=144.5
s=12.02... m
.
Ed